OpenStudy (anonymous):

There is a circle with two tangents that meet at point p, The lower tangent of the circle is 22cm. There is a 65 degree central angle formed by the radius and a line that goes from the center of the circle to the intersection of the two tangents. How long is the diameter?

OpenStudy (anonymous):

 \[by sine formula \frac{ r }{ \sin 25 }=\frac{ 25 }{ \sin 65 }\] calculate r, then d=2r

OpenStudy (anonymous):

Thank you, but why does this work?

OpenStudy (anonymous):

correction write 22/sin 65 in place of 25/sin 65